### v3

# Conversational Query Reformulation with the Guidance of Retrieved Documents

Abstract
Conversational search seeks to retrieve relevant passages for the given questions in conversational question answering. Conversational Query Reformulation (CQR) improves conversational search by refining the original queries into de-contextualized forms to resolve the issues in the original queries, such as omissions and coreferences. Previous CQR methods focus on imitating human written queries which may not always yield meaningful search results for the retriever. In this paper, we introduce GuideCQR, a framework that refines queries for CQR by leveraging key information from the initially retrieved documents. Specifically, GuideCQR extracts keywords and generates expected answers from the retrieved documents, then unifies them with the queries after filtering to add useful information that enhances the search process. Experimental results demonstrate that our proposed method achieves state-of-the-art performance across multiple datasets, outperforming previous CQR methods. Additionally, we show that GuideCQR can get additional performance gains in conversational search using various types of queries, even for queries written by humans. ***The code and pre-trained model will be publicly released upon the acceptance of the paper.
Conversational Query Reformulation with the Guidance of Retrieved Documents
Jeonghyun Park and Hwanhee Lee Department of Artificial Intelligence, Chung-Ang University, Seoul, Korea {tom0365, hwanheelee}@cau.ac.kr
1 Introduction
In a conversational question-answering task (ConvQA), conversational search aims to retrieve relevant passages that provide the necessary information to answer the current query. This process occurs within the framework of a multi-turn conversation, where each query builds upon the previous interactions. The questions in ConvQA often involve challenges like omissions and coreferences, which make it difficult to achieve the desired search results using the original query. Previous research focuses on transforming queries into stand-alone forms to understand their intent better, making them more independent and robust Mo et al. (2023a). This process, known as Conversational Query Reformulation (CQR), helps clarify the original queries and enhances query understanding.
With the advent of Large Language Models (LLMs), their application in CQR methods has become increasingly widespread. A recent study Jagerman et al. (2023) involves instructing an LLM to generate relevant passages related to the query. This method makes the query more reasonable and human-understandable by expanding the sentence. LLM4CS Mao et al. (2023) also involves using LLMs to rewrite the query through different prompting methods and aggregation techniques. However, although these approaches aim to create human-friendly and easily comprehensible queries, they may not always yield the effective results that retrievers desire. For instance, as shown at the bottom of Figure 1, although the reformulated query may seem less fluent due to the additional keywords, it achieves a higher retriever score than a base query because of its increased similarity to the ground-truth document. This example highlights the necessity of prioritizing retriever-optimized queries over human-friendly ones to enhance CQR methods. To achieve this, we observe that the initially retrieved documents from a retrieval process with a base query set can help the generation of retrieval-optimized queries. Specifically, we find that although some of the documents cannot provide clues for answering the questions, most of the retrieved documents contain words or signals that are highly influential in subsequent retrieval processes. For instance, in the example illustrated in Figure 1, the addition of the terms ’interchangeable’, an important keyword found in the gold passage within the guided documents, enhances the performance that the original query alone cannot achieve. In this way, signals in the initially retrieved documents obtained through a base query can guide the search for the gold passage.
In this paper, we propose GuideCQR, an effective CQR framework designed to generate retrieval-friendly conversational queries by leveraging the guidance of initially retrieved documents from the base query. GuideCQR first obtains the initially retrieved documents through a retrieval process using the baseline query reformulated by LLMs. We then perform a re-ranking process to refine the order of the retrieved documents based on their similarity to the query. Through this re-ranking process, we carefully select a small number of documents from the initially retrieved ones to extract signals that are more likely to contain the gold passage. Next, we extract the keywords and generate the expected answer from the re-ranked guided documents to obtain components for creating a retriever-friendly query. We then independently filter both the keywords and the expected answers based on their similarity to the original query and previous utterances to eliminate redundant information. Finally, we construct the final query set by concatenating the filtered keywords and expected answer with the baseline query.
Experimental results show that GuideCQR achieves state-of-the-art across several CQR datasets compared to the baseline systems. We also demonstrate the robustness of the GuideCQR framework in enhancing the retrievability of queries, making them more retriever-friendly. Furthermore, we validate the effectiveness of our method by analyzing the overlap between the augmented keywords and the relevant documents for the given question, confirming the alignment between the augmented queries and document content. Additionally, we show that the proposed method achieves notable improvements across various query sets, underscoring its adaptability.
2 Related Works
2.1 Conversational Search
The goal of conversational search Gao et al. (2022) is to retrieve passages containing the necessary information to answer the current query within a multi-turn conversation. To get the desired answers from the given conversational dialogue, understanding the meaning of the query is important in this field Mo et al. (2023b). However, queries in ConvQA have several problems. For instance, queries may include pronouns such as "he," "she," or "it," necessitating the identification of the referred entity, and there is also the challenge of omission, where essential information is not included Mao et al. (2023); Wang et al. (2023).
To address such challenges in conversational search, two main methods that enable conversational search are conversational dense retrieval and conversational query reformulation. Conversational dense retrieval focuses on enhancing the representation of the current query by incorporating historical context achieved through training dense retrievers Qu et al. (2020). On the other hand, conversational query reformulation transforms the conversational search into a traditional ad-hoc search by converting the entire search session into a single standalone query Elgohary et al. (2019).
In our work, we focus on utilizing the dialogue history to contextualize query rewrites. Additionally, we employ the multi-turn dialogue as a criterion for filtering keywords and answers, which are essential components of GuideCQR.
2.2 Conversational Query Reformulation
CQR methods aim to transform queries into de-contextualized forms, enabling them to be understood based solely on their content and ensuring they convey the intended meaning effectively. These methods are designed to enhance conversational search performance by refining and expanding user queries within a conversational context.
While previous approaches, such as human-rewritten queries Lin et al. (2020); Voskarides et al. (2020a); Yu et al. (2020), have attempted to improve query clarity, they often fall short in optimizing retrieval performance. Although human rewrites may enhance readability for users, they do not always align with the needs of retrieval systems.
To address these challenges, recent studies have introduced methods like CONQRR Wu et al. (2022), which directly optimizes query rewriting for retrieval through Self-Critical Sequence Training Rennie et al. (2017), and ConvGQRMo et al. (2023a), which improves retrieval performance by combining query rewriting with query expansion. Additionally, IterCQR Jang et al. (2024) iteratively trains the CQR model by using information retrieval (IR) signals as rewards. The integration of large language models (LLMs) in CQR has also shown promise in recent work Mao et al. (2023).
To the best of our knowledge, no prior study has directly utilized the signals within retrieved documents to apply CQR. Our proposed method, GuideCQR, focuses on transforming queries by leveraging information from the retrieved documents alongside the baseline query during the retrieval process.
3 GuideCQR
We propose GuideCQR, a novel framework designed to reformulate conversational queries by utilizing guidance from initially retrieved documents. Figure 2 provides an overview of the proposed query reformulation process. GuideCQR consists of three stages: We first retrieve an initial set of guided documents using a query reformulated by LLM. This step retrieves documents that are likely to contain gold passages to guide the query reformulation process (Sec 3.1). Next, we extract keywords and generate expected answers from the guided documents, creating components that contribute to making the query more retriever-friendly (Sec 3.2). Finally, we apply a filtering process that evaluates both keywords and answers based on their similarity scores relative to the baseline query and dialogue history. We then unify and concatenate these components back to the baseline query (Sec 3.3).
3.1 Guided Documents Retrieval
3.1.1 Initial Documents Retrieval
In the process of developing a retriever-friendly query, initially retrieved documents can play a crucial role as guiding resources by providing foundational insights for the CQR process. For example, in the contents of retrieved documents, these signals can include critical keywords or contextual signs that are necessary to the search the gold passages such as "cure," "rate," and "curable" from the document shown in Figure 2.
Inspired by these points, GuideCQR starts from getting initially retrieved documents to gain meaningful signals to the retriever. We obtain these initial documents by retrieving the documents using the baseline query set generated by LLM. We denote baseline query as ,
where represents an operation that resolves omissions or coreferences using OpenAI gpt3.5-turbo-16k and denotes the raw question in the dataset. denotes the dialogue history of , especially consisting of previous dialogues’ queries, rewritten queries and responses by human annotators. Based on both the and , we generate the . Using this , we obtain the initial documents composed of 2,000 documents for each sample ID, which are crucial for creating our final query set.
3.1.2 Re-ranking Documents
To further improve the quality of the guided set of documents, we implement a re-ranking process to generate initially retrieved documents. By reordering the initial documents using a different retrieval model, we aim to capture better documents by reducing biases that may arise from relying on a single retriever. Specifically, we employ Sentence-Transformer Reimers and Gurevych (2019) to re-rank the top 2,000 documents for each sample ID, selecting the final guided set by choosing the top 10 documents based on their similarity scores to the query.
3.2 Generating Retriever-friendly Query
Since many retrieved documents may contain influential words or signals that significantly impact subsequent retrieval stages, identifying key elements from these documents is crucial for constructing more effective queries. Based on the guided documents obtained from the previous step, we generate retriever-friendly queries with respect to two aspects: Augmenting Keywords, Expected Answer generation
3.2.1 Augmenting Keywords
Keywords play a critical role in forming retriever-friendly queries, as they capture the most relevant and significant terms within the document. To obtain these keywords, we utilize the guided documents.
We extract keywords by leveraging BERT Devlin et al. (2019) using the keyword extraction method as in KeyBERT Grootendorst (2020). We use BERT to compute embeddings, creating a representation for the entire document. Subsequently, we extract embeddings for N-gram words/phrases in the query. By computing cosine similarity, we determine which words/phrases are most similar to the document. We choose the keywords with the highest similarity.
In line with this principle, we augment the base queries with keywords through two key aspects. The first aspect concerns determining the number of documents to extract keywords from, which defines the level of guidance we aim to achieve. For each document, we extract keywords of a specified span length. The second aspect focuses on the span length of the keywords, which refers to the number of tokens, or the number of keywords, selected per document. For example, when augmenting top-2 span-3 keywords, we extract three keywords from each of the top two documents, resulting in a total of six keywords. Using this approach, we augment a total keyword list composed of keywords from the top- documents and span length :
where denotes unit keyword, is the number of documents and is keyword span length. For example, keywords such as "early", "throat", "cancer", "high", and "cure" can be crucial for the document as shown in Figure 2.
3.2.2 Expected Answer Generation
Expected answers for the given query can also serve as valuable signals to the query, providing direct insights into the user’s intent and thereby improving the relevance and accuracy of search results. We find that when generating the expected answer, rather than instructing language models to generate potential answers to the query, guided documents may contain gold answers to the query and can play a crucial role as a pool for finding the answers to reformulate query.
Hence, we also utilize the guided documents in generating expected answers as additional information to the query:
where represents answer list and denotes a unit answer extracted from a single document. To generate these answers, both the query and relevant context are necessary. We use the query as the baseline query and the context as the guided documents. For each document, we generate a single expected answer, such as "throat cancer has a high cure rate," as illustrated in Figure 2. We then aggregate these answers into a single sentence, treating the entire set as a unified response. Since each span contributes meaningfully to the final output, we avoid analyzing individual tokens individually. Consequently, this process produces k distinct answers, with one derived from each of the top-k documents.
3.3 Filtering and Unify
We observe that redundant elements, such as the keyword ’rate’ as shown in Figure 2, can emerge from both the augmented keywords and the generated answers. We explain this is because GuideCQR might augment the keywords or answers generated from irrelevant documents. So, irrelevant signals from irrelevant documents can have a negative impact when creating retriever-friendly queries.
To address this, we introduce an additional filtering stage to more effectively remove irrelevant keywords and answers from the reformulated query. This filtering process is guided by a metric we define as FilterScore, which leverages both and , calculated using cosine similarity, as follows:
where cosSim denotes cosine similarity ranging from 0 to 1. Based on , we firstly define as follows:
where query represents the embedding of the current turn’s query, and item refers to the embedding of either a keyword or an answer sentence. So is cosine similarity between query and item ranging from 0 to 10. And we define :
where history is the history query list of current utterances. So HistoryScore is the maximum cosine similarity value between the dialogue history query element in the history query list and current item. Finally, we define the FilterScore as the average of the QueryScore and the HistoryScore, ranging from 1 to 10:
History queries play a crucial role in understanding the current query, as in a dialogue, comprehending the question is more effective when based on the preceding conversation. Thus, we account for history queries through the , which allows us to traverse the entire dialogue and capture the global context from past to present. Using this , we can eliminate signals that are irrelevant to the current dialogue. Since keywords and answers may vary in retriever-friendliness across different datasets, we treat them as distinct units and apply different filtering scores rather than using the same score for both.
Using the , we filter out keywords and answers with a score below the specified threshold. Finally, we unify the remaining keywords and answers and integrate them into the to construct the final reformulated query as follows:
where and are remaining keywords, answers.
| CAsT-19 | CAsT-20 | QReCC | ||||||||
| Methods | MRR | NDCG@3 | R@10 | MRR | NDCG@3 | R@10 | MRR | NDCG@3 | R@10 | Avg |
| 41.2 | 24.3 | 5.8 | 23.2 | 15.4 | 5.5 | 10.2 | 9.3 | 15.7 | 16.7 | |
| Transformer++ | 69.6 | 44.1 | - | 29.6 | 18.5 | - | - | - | - | - |
| CQE-Sparse | 67.1 | 42.3 | - | 42.3 | 27.1 | - | 32.0 | 30.1 | 51.3 | - |
| QuReTeC | 68.9 | 43.0 | - | 43.0 | 28.7 | - | 35.0 | 32.6 | 55.0 | - |
| T5QR | 70.1 | 41.7 | - | 42.3 | 29.9 | - | 34.5 | 31.8 | 53.1 | - |
| ConvGQR | 70.8 | 43.4 | - | 46.5 | 33.1 | - | 42.0 | 39.1 | 63.5 | - |
| CONVINV | 74.2 | 44.9 | - | 47.6 | 33.4 | - | - | - | - | - |
| CHIQ | 73.3 | 50.5 | 12.9 | 54.0 | 38.0 | 19.3 | 47.0 | 44.2 | 70.8 | 45.5 |
| IterCQR | - | - | - | - | - | - | 42.9 | 40.2 | 65.5 | - |
| 72.2 | 45.2 | 12.2 | 53.4 | 37.2 | 18.2 | 43.8 | 51.9 | 66.4 | 44.5 | |
| LLM4CS | 77.6 | 51.5 | 10.9 | 61.5 | 45.5 | 17.0 | 44.8 | 42.1 | 66.4 | 46.3 |
| GuideCQR (ours) | 81.8 | 53.7 | 13.9 | 59.0 | 42.7 | 21.1 | 47.2 | 54.4 | 67.0 | 48.9 |
| Human Rewrite | 74.0 | 46.1 | 12.9 | 59.1 | 42.2 | 21.0 | 43.1 | 51.6 | 58.6 | 45.4 |
4 Experiments
4.1 Datasets and Metrics
We utilize three CQR benchmark datasets TREC CAsT-19 (Dalton et al., 2020), TREC CAsT-20 (Dalton et al., 2021) and QReCC dataset Anantha et al. (2021) for our work. CAsT-19 and CAsT-20 consist of 50 and 25 conversations, respectively. Unlike CAsT-19, CAsT-20 is more realistic and complex because its queries are based on information needs drawn from commercial search logs, and they can reference prior system responses. CAsT series datasets assign a relevance score ranging from 1 to 4 to each passage, indicating the degree of relevance to the query, with documents scoring a 4 being considered gold passages. For the QReCC dataset, each query is paired with a single gold passage different from CAsT dataset. QReCC dataset includes a training set and a test set, and we sample 2K conversations from the training set to create a development set, following the previous work Kim and Kim (2022).
Following the prior CQR research (Mo et al., 2023a; Mao et al., 2023; Jang et al., 2024), we use three widely used evaluation metrics for CQR to compare the performance: Mean Reciprocal Rank (MRR), Normalized Discounted Cumulative Gain at three documents (NDCG@3), Recall@10. We utilize pytrec_eval (Van Gysel and de Rijke, 2018) tool to compute the score.
4.2 Baselines
We compare GuideCQR with the following CQR methods: (1) Transformer++ (Vakulenko et al., 2021a): A GPT-2 based CQR model fine-tuned on CANARD Elgohary et al. (2019) dataset. (2) CQE-Sparse (Lin et al., 2021b): A weakly-supervised method to select important tokens only from the context via contextualized query embeddings. (3) QuReTeC (Voskarides et al., 2020b): A weakly-supervised method to train a sequence tagger to decide whether each term contained in a historical context should be added to the current query. (4) T5QR (Lin et al., 2020): A conversational query rewriter based on T5, trained using human-generated rewrites. (5) ConvGQR (Mo et al., 2023a): A query reformulation framework that combines query rewriting with generative query expansion. (6) CONVINV (Cheng et al., 2024): Framework that transforms conversational session embeddings into interpretable text using Vec2Text. (7) CHIQ Mo et al. (2024): A two-step method that leverages the capabilities of LLMs to resolve ambiguities in the conversation history before query rewriting. (8) IterCQR Jang et al. (2024): Improves a CQR framework through iterative training method based on the similarity between the passage and the query. (9) LLM4CS: Query rewriting based on LLM and various prompting methods. We build GuideCQR upon this prompted QR method using (REW + Maxprob) and evaluate using self-consistency.
4.3 Implementation Details
To generate a baseline query for CAsT-19 and CAsT-20, we utilize OpenAI gpt3.5-turbo-16k combined with the Maxprob approach as proposed in LLM4CS. We simply generate this query by instructing LLM to rewrite the raw query based on dialog history and sampling the highest generation probability with LLMs, resolving only omissions and coreferences. For the QReCC dataset, rather than generating ourselves, we adopt the final query output generated by InfoCQR Ye et al. (2023) as . This approach leverages prompts from gpt3.5-turbo to enhance query generation.
4.4 Performance Comparison
Main Results
As shown in Table 1, GuideCQR significantly enhances performance metrics compared to the across all datasets. GuideCQR achieves state-of-the-art performance in terms of average score. In addition, GuideCQR achieves either the best or second-best performance compared to all baselines and demonstrates its robustness. Specifically, GuideCQR outperforms LLM4CS by 5.4% in MRR on the CAsT-19 dataset and by 29.2% in NDCG@3 on the QReCC dataset. For CAsT-20, still its performance is almost the second-best. These results highlight the robustness and effectiveness of GuideCQR.
We attribute the lower score for CAsT-20 compared to CAsT-19 to the increased complexity of the topics in CAsT-20. Specifically, in the retrieval process, we set the parameter rel threshold which represents the minimum query relevance score for a document to be considered relevant to the query. CAsT-19 uses this threshold as 1 and CAsT-20 uses 2, so the minimum criteria for relevance is higher in CAsT-20. As a result, the number of relevant documents in CAsT-20 is significantly lower than in CAsT-19. Furthermore, the query relevance score file for CAsT-20 contains relatively few documents with high relevance scores; in most cases, the score is 0. This makes CAsT-20 a more challenging dataset compared to CAsT-19 for applying GuideCQR. In other words, the signals derived from irrelevant guiding documents may not perform effectively to GuideCQR in CAsT-20.
Ablation Study
To evaluate the effectiveness of individual components involved in creating retriever-friendly queries in our proposed CQR framework, we conduct ablation study. As demonstrated in Table 2, removing any part of GuideCQR leads to a decrease in performance, indicating that each component plays an important role in GuideCQR.
| CAsT-19 | CAsT-20 | |||
| MRR | NDCG@3 | MRR | NDCG@3 | |
| GuideCQR (ours) | 81.8 | 53.7 | 59.0 | 42.7 |
| - w/o keywords | 72.0 | 47.3 | 54.7 | 38.6 |
| - w/o answers | 81.0 | 52.2 | 57.9 | 41.6 |
| - w/o filtering | 79.7 | 52.2 | 58.4 | 43.2 |
4.5 Analysis
Performance among Amount of Initial Guided documents
We evaluate the impact and robustness of each step in the GuideCQR setup. Initially, we adjust the number of guided documents to observe the proper quantity and present the results in Table 3. Our findings demonstrate that increasing the number of guided documents consistently enhances performance. However, retrieving an excessive number of documents leads to longer inference times. Therefore, we limit the number of initial documents to 2,000 for our main experiment.
| # of initial guided documents | MRR | NDCG@3 | Avg |
| 10 | 75.7 | 50.5 | 63.1 |
| 100 | 78.2 | 51.4 | 64.8 |
| 1,000 | 79.5 | 52.1 | 65.8 |
| 2,000 | 81.0 | 52.2 | 66.6 |
Performance among Top-k Documents for Keyword and Answer
As presented in Table 4, we examine the impact of varying the number of documents used in augmenting keywords and generating expected answers. For instance, when extracting keywords and answers from the top two documents, we generate one keyword-answer pair per document, resulting in two pairs in total. Our findings indicate that incorporating a larger number of documents generally provides additional information, which can improve performance. However, using too many documents may cause a decline in overall performance similar to the amount of initial guided documents.
| Keyword | Answer | |||||
| MRR | NDCG@3 | Avg | MRR | NDCG@3 | Avg | |
| 72.2 | 45.2 | 58.7 | 72.2 | 45.2 | 58.7 | |
| top-2 | 78.3 | 50.1 | 64.2 | 76.7 | 48.3 | 62.5 |
| top-3 | 80.4 | 52.1 | 66.2 | 77.5 | 49.0 | 63.2 |
| top-4 | 81.0 | 52.2 | 66.6 | 76.9 | 50.0 | 63.4 |
| top-5 | 80.2 | 52.3 | 66.2 | 77.5 | 49.6 | 63.5 |
Performance among Keyword Span Length
We also investigate the impact of varying keyword spans in augmenting the keyword stage and present the result in Table 5. We observe that increasing the keyword span generally enhances performance to a certain extent. However, we find that too many signals can degrade performance due to inclusion of the redundant information in the query.
| CAsT-19 | CAsT-20 | |||
| MRR | NDCG@3 | MRR | NDCG@3 | |
| 72.2 | 45.2 | 53.4 | 37.2 | |
| 5 span | 76.9 | 49.9 | 57.7 | 41.5 |
| 10 span | 78.8 | 52.0 | 53.7 | 39.8 |
| 15 span | 81.0 | 52.2 | 54.1 | 40.0 |
| 20 span | 80.1 | 52.0 | 54.6 | 39.8 |
Query Relevance Analysis
To demonstrate the effectiveness of the keyword augmentation step, we evaluate the precision score using the following formula:
where represents the number of unique matched keywords found in the guided documents that have query relevance score of 4 and represents the total number of unique augmented keywords. Note that the document that has a relevance score of 4 is regarded as the most relevant document with user queries in CAsT datasets. Matching keywords in these documents can confirm that augmenting keywords plays a crucial role in generating retriever-friendly queries. As shown in Table 6, a longer keyword span results in a higher precision score, highlighting the significance of keyword augmentation. However, an excessive number of documents may negatively impact the overall score.
| Span Length | Precision |
| 5 | 0.6684 |
| 10 | 0.7018 |
| 15 | 0.7024 |
| 20 | 0.7045 |
Failure Case Study
We analyze a representative failure case of our proposed method by sampling the final query in the CAsT-19 dataset. Following previous works Mao et al. (2023), we first set the relevance threshold for CAsT-19 at 1, meaning that we consider retrieved documents with a score of 1 or higher relevant to the query. Therefore, in the sampling failure case process, if the top-ranked (first position) retrieved document has a relevance score of 0, we classify it as a failure case. As shown in Figure 3, when retrieving documents using , the relevance score drops from 2 to 0 compared to , indicating a shift from relevant to irrelevant content. In , the query pertains to the main character of The NeverEnding Story, but the red-marked keywords and blue-marked answers in the retrieved documents refer to actual authors or actors, leading to a failure case. This observation suggests that performance tends to degrade when unrelated keywords and answers are introduced into the query, deviating from the original context of .
Adaptability of GuideCQR
GuideCQR is a highly effective framework as it operates as a query expansion method, making it adaptable for use with other methods or queries simultaneously. To demonstrate its effectiveness, we conduct experiments using both human-rewritten queries and from the CAsT-19 dataset. The refers to the original query without any reformulation applied. As shown in Table 7, both and human-rewritten queries show performance improvements. This demonstrates that our framework can be applied to any query or CQR method to achieve enhanced performance.
| MRR | NDCG@3 | |
| 41.2 | 24.3 | |
| + GuideCQR | 47.0(+5.8) | 28.8(+4.5) |
| Human Rewrite | 74.0 | 46.1 |
| + GuideCQR | 82.8(+8.8) | 53.0(+6.9) |
5 Conclusion
We present the GuideCQR, a novel query reformulation framework for conversational search. GuideCQR effectively reformulates conversational queries to be more retriever-friendly by leveraging guided documents. GuideCQR not only better aligns queries with user intent but also significantly enhances both the precision and relevance of retrieval results, offering a meaningful advancement in conversational search systems. Experimental results across various datasets and metrics confirm the capability of GuideCQR over human rewritten query and previous CQR methods.
Limitations
Our proposed framework generally takes more time to reformulate conversational queries compared to conventional CQR methods. In particular, the stages involving keyword augmentation and expected answer generation tend to be more time-consuming in GuideCQR, especially as the keyword span length and the number of documents increase. Future research will need to address these inefficiencies to reduce the overall processing time.
Moreover, the optimal keyword span length and number of documents may vary depending on the dataset. While incorporating more signals can enhance search results, the ideal amount of signals across diverse datasets remains undetermined. Thus, optimizing signal usage for different datasets represents a promising avenue for future exploration.
Despite these limitations, we believe our framework offers a novel perspective on addressing CQR challenges. As a result, we expect it will positively influence the future development of CQR methods.
Ethics Statement
This research leverages multiple publicly available datasets for conversational query reformulation. We have accurately cited all the papers and sources used in our study. We intend to publish the final query and the code including the pre-trained model for our proposed framework once the paper is accepted.
Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
Appendix A Implementation Details
To re-rank the guided documents as in section 3.1.2, we select the ember-v1 †††https://huggingface.co/llmrails/ember-v1 and mxbai-embed-large-v1 ‡‡‡https://huggingface.co/mixedbread-ai/mxbai-embed-large-v1 models as the first and second re-ranking Sentence-Transformer models, respectively. And we use OpenAI text-embedding-3-small for generating embeddings for keywords and answers in the filtering stage in GuideCQR. We use different embedding models at each stage to mitigate potential biases caused by a single model. We use KeyBERT (Grootendorst, 2020) for augmenting keywords. And we use RoBERTa-Base-Squad2 §§§https://huggingface.co/deepset/roberta-base-squad2 for generating the expected answer. We use the most common query among the five candidate queries provided by the baseline Mao et al. (2023) during the re-ranking, answer generation and filtering process in CAsT datasets. This is because Mao et al. (2023) provides a set of five candidate queries when generating their queries, so we make representative among them. And in augmenting keyword process, the number of keywords extracted can be fewer than the span length because the number of keywords reasonable in each document can be less than the length of the keyword span. This means that the keyword span length is maximum number of keywords for each document.
We conduct our experiments using an AMD EPYC 7313 CPU (3.0 GHz) paired with four NVIDIA RTX 4090 GPUs. We use Python 3.11.5 and PyTorch 2.3.1 for the software environment .
Appendix B Query Analysis
We present example queries reformulated by our method and by humans in Table 10. While our final query, , may not appear intuitive from a human perspective, it yields better search results during the retrieval stage.
Appendix C Selecting FilterScore
We experiment with various settings to find the optimal hyperparameter with the validation set for the proposed method. Before filtering, we use top-4 span 15 keywords and top-10 answers for CAsT-19, and top-5 span 5 keywords and top-10 answers for CAsT-20 and top-1 span 10 keywords and top-10 answer for QReCC. And GuideCQR uses these keywords and answers as our final query set. We select the for each keyword and answer by experimenting with validation sets for the datasets to find the optimal scores. We achieve our final results using a of (1, 1.9) for each keyword and answer in CAsT-19, (0.1, 1.95) in CAsT-20, and (0.5, 9) in QReCC.
Appendix D Process of generating of CAsT
We generate of CAsT datasets by utilizing the LLM4CS REW GitHub code which employs GPT-3.5-turbo-16k for prompting Mao et al. (2023). As illustrated in Table 11, the prompt consists of three parts: instruction, demonstration, tail instruction. The final prompt is a concatenation of these three components. Using this prompt, GuideCQR generate by refining with , resolving any coreferences or omissions.
Appendix E Dialog history usage in GuideCQR
In conversational search, utilizing dialog history is crucial. GuideCQR leverages the conversation history in two key ways. First, we use dialog history when generating , as it helps refine the current query by resolving any ambiguities, coreferences, or omissions from previous interactions. Additionally, we incorporate history in filtering augmented keywords and expected answers. By measuring the similarity with the history, we can select the most relevant keywords or answers. Through these methods, GuideCQR effectively integrates history, demonstrating its capability as a conversational search system.
Appendix F Dataset Statistics
CAsT
Table 8 presents statistics for the CAsT datasets, including the number of conversations, queries, and documents. Both CAsT datasets share the same document collection and provide passage-level relevance judgments, as well as human rewrites for each turn.
| CAsT-19 | CAsT-20 | |
| # Conv | 50 | 25 |
| # Questions | 479 | 216 |
| # Documents | 38M | 38M |
QReCC
Table 9 shows the statistics for the QReCC dataset, which contains approximately 54M documents. We use a final test set of 8,209 questions, following Ye et al. (2023), after removing invalid gold passage labels. This test set includes 6,396 questions for QuAC-Conv, 1,442 for NQ-Conv, and 371 for TREC-Conv.
| Train | Dev | Test | |
| QReCC | |||
| # Conv | 8,823 | 2,000 | 2,775 |
| # Questions | 51,928 | 11,573 | 16,451 |
| QuAC-Conv | |||
| # Conv | 6,008 | 1,300 | 1,816 |
| # Questions | 41,395 | 8,965 | 12,389 |
| NQ-Conv | |||
| # Conv | 2,815 | 700 | 879 |
| # Questions | 10,533 | 2,608 | 3,314 |
| TREC-Conv | |||
| # Conv | 0 | 0 | 80 |
| # Questions | 0 | 0 | 748 |
| RawQuery | What are its other competitors? |
| Human-rewritten | What are Netflix’s other competitors than Blockbuster? |
| What are Netflix’s other competitors? | |
| What are Netflix’s other competitors? netflix competitors subscription streaming amazon instant hulu plus amazon netflix competes viaplay hbo nordic cmore asia netflix competes hooq sky demand competitors apple itunes amazon cinemanow hulu lovefilm google netflix fandangonow ultraflix voddler competitors Amazon Instant Video Service and the Hulu Plus service Presto, Stan and Quickflix FandangoNow, UltraFlix, and Voddler : Competitors brick and mortar video rental stores |
| Instruction | For an information-seeking dialog, please help reformulate the question into a rewrite that can fully express the user’s information needs without the need for context. |
| Demonstration | I will give you several example multi-turn dialogs, where each turn contains a question, a response, and a rewrite that you need to generate. |
| Tail Instruction | Now, you should give me the rewrite of the **Current Question** under the **Context**. The output format should always be: Rewrite: $Rewrite. Note that you should always try to rewrite it. Never ask for clarification or say you don’t understand it in the generated rewrite. Go ahead! |

### v4

@@
 Additionally, we show that GuideCQR can get additional performance gains in conversational search using various types of queries, even for queries written by humans.
-***The code and pre-trained model will be publicly released upon the acceptance of the paper.
+111The code and pre-trained model will be publicly released upon the acceptance of the paper.
 Conversational Query Reformulation with the Guidance of Retrieved Documents
@@
 In this paper, we propose GuideCQR, an effective CQR framework designed to generate retrieval-friendly conversational queries by leveraging the guidance of initially retrieved documents from the base query.
-GuideCQR first obtains the initially retrieved documents through a retrieval process using the baseline query reformulated by LLMs.
+GuideCQR first obtains the initially retrieved documents through a retrieval process using the baseline query reformulated by LLM.
 We then perform a re-ranking process to refine the order of the retrieved documents based on their similarity to the query.
@@
 We then independently filter both the keywords and the expected answers based on their similarity to the original query and previous utterances to eliminate redundant information.
-Finally, we construct the final query set by concatenating the filtered keywords and expected answer with the baseline query.
+Finally, we construct the final query set by concatenating the filtered keywords and expected answers with the baseline query.
 Experimental results show that GuideCQR achieves state-of-the-art across several CQR datasets compared to the baseline systems.
 We also demonstrate the robustness of the GuideCQR framework in enhancing the retrievability of queries, making them more retriever-friendly.
-Furthermore, we validate the effectiveness of our method by analyzing the overlap between the augmented keywords and the relevant documents for the given question, confirming the alignment between the augmented queries and document content.
+Furthermore, we validate the effectiveness of our method by analyzing the overlap between the augmented keywords and the relevant documents for the given question.
 Additionally, we show that the proposed method achieves notable improvements across various query sets, underscoring its adaptability.
@@
 (2023).
-To address such challenges in conversational search, two main methods that enable conversational search are conversational dense retrieval and conversational query reformulation.
-Conversational dense retrieval focuses on enhancing the representation of the current query by incorporating historical context achieved through training dense retrievers Qu et al.
+To address such challenges in conversational search, two main methods that enable conversational search are Conversational Dense Retrieval (CDR) and CQR.
+CDR focuses on enhancing the representation of the current query by incorporating historical context achieved through training dense retrievers Qu et al.
 (2020).
-On the other hand, conversational query reformulation transforms the conversational search into a traditional ad-hoc search by converting the entire search session into a single standalone query Elgohary et al.
+On the other hand, CQR transforms the conversational search into a traditional ad-hoc search by converting the entire search session into a single standalone query Elgohary et al.
 (2019).
@@
 (2024) iteratively trains the CQR model by using information retrieval (IR) signals as rewards.
-The integration of large language models (LLMs) in CQR has also shown promise in recent work Mao et al.
+The integration of LLMs in CQR has also shown promise in recent work Mao et al.
 (2023).
@@
 Finally, we apply a filtering process that evaluates both keywords and answers based on their similarity scores relative to the baseline query and dialogue history.
-We then unify and concatenate these components back to the baseline query (Sec 3.3).
+We then unify and concatenate these components with the baseline query to construct the final query for post retrieval (Sec 3.3).
 3.1 Guided Documents Retrieval
@@
 We denote baseline query as ,
-where represents an operation that resolves omissions or coreferences using OpenAI gpt3.5-turbo-16k and denotes the raw question in the dataset.
+where represents an operation that resolves omissions or coreferences using OpenAI gpt3.5-turbo-16k and denotes the raw question in the dataset, without any reformulation applied.
 denotes the dialogue history of , especially consisting of previous dialogues’ queries, rewritten queries and responses by human annotators.
 Based on both the and , we generate the .
-Using this , we obtain the initial documents composed of 2,000 documents for each sample ID, which are crucial for creating our final query set.
+Using this , we obtain the initial documents composed of 2,000 documents for each question.
+They are crucial for creating our final query set and strong foundation for identifying meaningful signals and guiding the retriever towards the most relevant passages.
 3.1.2 Re-ranking Documents
@@
 By reordering the initial documents using a different retrieval model, we aim to capture better documents by reducing biases that may arise from relying on a single retriever.
-Specifically, we employ Sentence-Transformer Reimers and Gurevych (2019) to re-rank the top 2,000 documents for each sample ID, selecting the final guided set by choosing the top 10 documents based on their similarity scores to the query.
+Specifically, we employ Sentence-Transformer Reimers and Gurevych (2019) to re-rank the top 2,000 documents for each question, selecting the final guided set by choosing the top 10 documents based on their similarity scores to the query.
 3.2 Generating Retriever-friendly Query
 Since many retrieved documents may contain influential words or signals that significantly impact subsequent retrieval stages, identifying key elements from these documents is crucial for constructing more effective queries.
-Based on the guided documents obtained from the previous step, we generate retriever-friendly queries with respect to two aspects: Augmenting Keywords, Expected Answer generation
+Based on the guided documents obtained from the previous step, we generate retriever-friendly queries by incorporating extra information to the query from two approaches: Augmenting Keywords, Expected Answer generation
 3.2.1 Augmenting Keywords
-Keywords play a critical role in forming retriever-friendly queries, as they capture the most relevant and significant terms within the document.
-To obtain these keywords, we utilize the guided documents.
-We extract keywords by leveraging BERT Devlin et al.
-(2019) using the keyword extraction method as in KeyBERT Grootendorst (2020).
-We use BERT to compute embeddings, creating a representation for the entire document.
-Subsequently, we extract embeddings for N-gram words/phrases in the query.
-By computing cosine similarity, we determine which words/phrases are most similar to the document.
-We choose the keywords with the highest similarity.
-In line with this principle, we augment the base queries with keywords through two key aspects.
-The first aspect concerns determining the number of documents to extract keywords from, which defines the level of guidance we aim to achieve.
-For each document, we extract keywords of a specified span length.
-The second aspect focuses on the span length of the keywords, which refers to the number of tokens, or the number of keywords, selected per document.
-For example, when augmenting top-2 span-3 keywords, we extract three keywords from each of the top two documents, resulting in a total of six keywords.
-Using this approach, we augment a total keyword list composed of keywords from the top- documents and span length :
+We find that keywords from the initial documents play a critical role in forming retriever-friendly queries, as they capture the most relevant and significant terms within the document.
+For example, keywords such as "early", "throat", "cancer", "high", and "cure" can be beneficial if they are augmented to the search query as shown in Figure 2.
+We leverage KeyBERT Grootendorst (2020) to extract keywords from the re-ranked documents.
+The process begins by using BERT Devlin et al.
+(2019) to compute embeddings of documents, which create a representation for the entire document.
+Embeddings for N-gram words or phrases within the document are then extracted.
+By calculating the cosine similarity between these words/phrases and the document representation, the method identifies which words/phrases are most similar to the document.
+In line with this principle, we enhance the base queries by augmenting keywords through two hyperparameters.
+The first one involves determining the number of documents to extract keywords from, which establishes the level of guidance we intend to provide.
+For each guided document, we extract keywords of a specified span length.
+The second aspect pertains to the span length, which indicates the number of tokens or keywords selected per document.
+For instance, when augmenting with top-2 span-3 keywords, we extract three keywords from each of the top two documents, yielding a total of six keywords.
+Consequently, we augment a total keyword list composed of keywords from the top- documents and span length :
 where denotes unit keyword, is the number of documents and is keyword span length.
-For example, keywords such as "early", "throat", "cancer", "high", and "cure" can be crucial for the document as shown in Figure 2.
 3.2.2 Expected Answer Generation
-Expected answers for the given query can also serve as valuable signals to the query, providing direct insights into the user’s intent and thereby improving the relevance and accuracy of search results.
-We find that when generating the expected answer, rather than instructing language models to generate potential answers to the query, guided documents may contain gold answers to the query and can play a crucial role as a pool for finding the answers to reformulate query.
-Hence, we also utilize the guided documents in generating expected answers as additional information to the query:
+Guided documents often include gold answers to the query, serving as a valuable resource for efficiently reformulating the query.
+Based on this idea, we use guided documents as context to generate expected answers to enhance the query.
+As illustrated in Figure 2, although it’s not obtained from the gold passage, expected answer is a concise and informative response that can potentially address the user’s query since they can provide direct insights into the user’s intent, thereby improving the relevance and accuracy of search results.
+Specifically, we generate these expected answers as follows:
 where represents answer list and denotes a unit answer extracted from a single document.
@@
 We use the query as the baseline query and the context as the guided documents.
-For each document, we generate a single expected answer, such as "throat cancer has a high cure rate," as illustrated in Figure 2.
-We then aggregate these answers into a single sentence, treating the entire set as a unified response.
-Since each span contributes meaningfully to the final output, we avoid analyzing individual tokens individually.
+We generate a single expected answer for each document.
 Consequently, this process produces k distinct answers, with one derived from each of the top-k documents.
@@
 We observe that redundant elements, such as the keyword ’rate’ as shown in Figure 2, can emerge from both the augmented keywords and the generated answers.
-We explain this is because GuideCQR might augment the keywords or answers generated from irrelevant documents.
-So, irrelevant signals from irrelevant documents can have a negative impact when creating retriever-friendly queries.
+We find that this is because GuideCQR might augment keywords or answers derived from irrelevant documents.
+Hence, irrelevant signals from irrelevant documents can have a negative impact when creating retriever-friendly queries.
 To address this, we introduce an additional filtering stage to more effectively remove irrelevant keywords and answers from the reformulated query.
-This filtering process is guided by a metric we define as FilterScore, which leverages both and , calculated using cosine similarity, as follows:
+We guide this filtering process through the metric FilterScore, which leverages both and , calculated using cosine similarity, as follows:
 where cosSim denotes cosine similarity ranging from 0 to 1. Based on , we firstly define as follows:
@@
 | ConvGQR | 70.8 | 43.4 | - | 46.5 | 33.1 | - | 42.0 | 39.1 | 63.5 | - |
-| CONVINV | 74.2 | 44.9 | - | 47.6 | 33.4 | - | - | - | - | - |
+| CONVINV | 74.2 | 44.9 | - | 47.6 | 34.4 | - | - | - | - | - |
 | CHIQ | 73.3 | 50.5 | 12.9 | 54.0 | 38.0 | 19.3 | 47.0 | 44.2 | 70.8 | 45.5 |
@@
 CAsT-19 and CAsT-20 consist of 50 and 25 conversations, respectively.
+Both CAsT datasets share the same document collection and provide passage-level relevance judgments, as well as human rewrites for each turn.
 Unlike CAsT-19, CAsT-20 is more realistic and complex because its queries are based on information needs drawn from commercial search logs, and they can reference prior system responses.
@@
 We utilize pytrec_eval (Van Gysel and de Rijke, 2018) tool to compute the score.
+4.2 Implementation Details
+To generate a baseline query for CAsT-19 and CAsT-20, we utilize OpenAI gpt3.5-turbo-16k combined with the Maxprob approach as proposed in LLM4CS.
+We simply generate this query by instructing LLM to rewrite the raw query based on dialog history and sampling the highest generation probability with LLMs, resolving only omissions and coreferences.
+For the QReCC dataset, rather than generating ourselves, we adopt the final query output generated by InfoCQR Ye et al.
+(2023) as .
+This approach leverages prompts from gpt3.5-turbo to enhance query generation.
 4.3 Baselines
@@
 (8) IterCQR Jang et al.
-(2024): Improves a CQR framework through iterative training method based on the similarity between the passage and the query.
+(2024): CQR framework through iterative refinement based on the similarity between the passage and the query.
 (9) LLM4CS: Query rewriting based on LLM and various prompting methods.
-We build GuideCQR upon this prompted QR method using (REW + Maxprob) and evaluate using self-consistency.
-4.3 Implementation Details
-To generate a baseline query for CAsT-19 and CAsT-20, we utilize OpenAI gpt3.5-turbo-16k combined with the Maxprob approach as proposed in LLM4CS.
-We simply generate this query by instructing LLM to rewrite the raw query based on dialog history and sampling the highest generation probability with LLMs, resolving only omissions and coreferences.
-For the QReCC dataset, rather than generating ourselves, we adopt the final query output generated by InfoCQR Ye et al.
-(2023) as .
-This approach leverages prompts from gpt3.5-turbo to enhance query generation.
 4.4 Performance Comparison
@@
 4.5 Analysis
-Performance among Amount of Initial Guided documents
-We evaluate the impact and robustness of each step in the GuideCQR setup.
-Initially, we adjust the number of guided documents to observe the proper quantity and present the results in Table 3.
-Our findings demonstrate that increasing the number of guided documents consistently enhances performance.
-However, retrieving an excessive number of documents leads to longer inference times.
-Therefore, we limit the number of initial documents to 2,000 for our main experiment.
-| # of initial guided documents | MRR | NDCG@3 | Avg |
-| 10 | 75.7 | 50.5 | 63.1 |
-| 100 | 78.2 | 51.4 | 64.8 |
-| 1,000 | 79.5 | 52.1 | 65.8 |
-| 2,000 | 81.0 | 52.2 | 66.6 |
+CAsT dataset includes multiple relevant passages, each with a relevance score. We focus on CAsT due to these unique label features.
 Performance among Top-k Documents for Keyword and Answer
-As presented in Table 4, we examine the impact of varying the number of documents used in augmenting keywords and generating expected answers.
-For instance, when extracting keywords and answers from the top two documents, we generate one keyword-answer pair per document, resulting in two pairs in total.
+As shown in Table 3, we verify the impact of varying the number of documents used in augmenting keywords and generating expected answers.
 Our findings indicate that incorporating a larger number of documents generally provides additional information, which can improve performance.
@@
 | top-5 | 80.2 | 52.3 | 66.2 | 77.5 | 49.6 | 63.5 |
-Performance among Keyword Span Length
-We also investigate the impact of varying keyword spans in augmenting the keyword stage and present the result in Table 5.
-We observe that increasing the keyword span generally enhances performance to a certain extent.
-However, we find that too many signals can degrade performance due to inclusion of the redundant information in the query.
-| CAsT-19 | CAsT-20 | |||
-| MRR | NDCG@3 | MRR | NDCG@3 | |
-| 72.2 | 45.2 | 53.4 | 37.2 | |
-| 5 span | 76.9 | 49.9 | 57.7 | 41.5 |
-| 10 span | 78.8 | 52.0 | 53.7 | 39.8 |
-| 15 span | 81.0 | 52.2 | 54.1 | 40.0 |
-| 20 span | 80.1 | 52.0 | 54.6 | 39.8 |
+Performance among Amount of Initial Guided documents
+| # of initial guided documents | MRR | NDCG@3 | Avg |
+| 10 | 75.7 | 50.5 | 63.1 |
+| 100 | 78.2 | 51.4 | 64.8 |
+| 1,000 | 79.5 | 52.1 | 65.8 |
+| 2,000 | 81.0 | 52.2 | 66.6 |
+We evaluate the impact and robustness of each step in the GuideCQR setup.
+Initially, we adjust the number of guided documents to observe the proper quantity and present the results in Table 4.
+Our findings demonstrate that increasing the number of guided documents consistently enhances performance.
+However, retrieving an excessive number of documents leads to longer inference times.
+Hence, considering the computation cost, we decide to use 2000 documents which is the point on the dev set where there is no additional performance improvement in at least one metric.
+We also use the same criteria for keyword span length.
+GuideCQR MRR NDCG@3 R@10 MRR NDCG@3 R@10 RawQuery 41.2 24.3 5.8 47.0(+5.8) 28.8(+4.5) 6.6(+0.8) Human Rewrite 74.0 46.1 12.9 82.8(+8.8) 53.0(+6.9) 14.4(+1.5) GPT3.5-turbo-16k 72.2 45.2 12.2 81.8(+9.6) 53.7(+8.5) 13.9(+1.7) Qwen2.5-7B-Instruct 62.9 39.3 10.2 65.0(+2.1) 42.3(+3.0) 11.7(+1.5) Mistral-7B-Instruct-v0.3 69.1 42.8 11.6 69.8(+0.7) 46.0(+3.2) 12.0(+0.4) EXAONE-3.5-7.8B-Instruct 66.4 41.5 11.1 72.6(+6.2) 46.7(+5.2) 12.5(+1.4)
 Query Relevance Analysis
 To demonstrate the effectiveness of the keyword augmentation step, we evaluate the precision score using the following formula:
-where represents the number of unique matched keywords found in the guided documents that have query relevance score of 4 and represents the total number of unique augmented keywords.
-Note that the document that has a relevance score of 4 is regarded as the most relevant document with user queries in CAsT datasets.
+where represents the number of unique matched keywords found in the guided documents that have query relevance score of and represents the total number of unique augmented keywords.
+Note that the document that has a relevance score of 1 or is regarded as the most relevant document with user queries in CAsT datasets.
 Matching keywords in these documents can confirm that augmenting keywords plays a crucial role in generating retriever-friendly queries.
-As shown in Table 6, a longer keyword span results in a higher precision score, highlighting the significance of keyword augmentation.
-However, an excessive number of documents may negatively impact the overall score.
-| Span Length | Precision |
-| 5 | 0.6684 |
-| 10 | 0.7018 |
-| 15 | 0.7024 |
-| 20 | 0.7045 |
+As shown in Table 6, the precision score for relevant passages is higher than that for irrelevant passages, indicating that augmenting keywords play a crucial role in retrieving relevant documents.
+| Precision | ||
+| Irrelevant | Passage (0) | 57.21 |
+| Relevant | Passage (1) | 67.39 |
+| Passage (2) | 71.36 | |
+| Passage (3) | 72.26 | |
+| Passage (4) | 70.45 |
 Failure Case Study
@@
 GuideCQR is a highly effective framework as it operates as a query expansion method, making it adaptable for use with other methods or queries simultaneously.
-To demonstrate its effectiveness, we conduct experiments using both human-rewritten queries and from the CAsT-19 dataset.
-The refers to the original query without any reformulation applied.
-As shown in Table 7, both and human-rewritten queries show performance improvements.
+To demonstrate its effectiveness, we conduct experiments using human-rewritten queries, and with different generator models from the CAsT-19 dataset.
+We create using GPT-3.5-turbo-16k to derive the final performance of GuideCQR.
+Furthermore, we conduct experiments with EXAONE-3.5-7.8B-Instruct Research (2024), Mistral-7B-Instruct-v0.3 Jiang et al.
+(2023), and Qwen2.5-7B-Instruct Team (2024).
+As shown in Table 5, , human-rewritten queries and with different generators show performance improvements.
 This demonstrates that our framework can be applied to any query or CQR method to achieve enhanced performance.
-| MRR | NDCG@3 | |
-| 41.2 | 24.3 | |
-| + GuideCQR | 47.0(+5.8) | 28.8(+4.5) |
-| Human Rewrite | 74.0 | 46.1 |
-| + GuideCQR | 82.8(+8.8) | 53.0(+6.9) |
 5 Conclusion
-We present the GuideCQR, a novel query reformulation framework for conversational search.
-GuideCQR effectively reformulates conversational queries to be more retriever-friendly by leveraging guided documents.
-GuideCQR not only better aligns queries with user intent but also significantly enhances both the precision and relevance of retrieval results, offering a meaningful advancement in conversational search systems.
+We present GuideCQR, a novel query reformulation framework utilizing initially retrieved documents from query set for conversational search.
+GuideCQR effectively reformulates conversational queries to be more retriever-friendly by extracting meaningful information from these guided documents.
 Experimental results across various datasets and metrics confirm the capability of GuideCQR over human rewritten query and previous CQR methods.
+We also show that GuideCQR can be effectively adapted to various types of re-written queries.
 Limitations
@@
 Appendix A Implementation Details
-To re-rank the guided documents as in section 3.1.2, we select the ember-v1 †††https://huggingface.co/llmrails/ember-v1 and mxbai-embed-large-v1 ‡‡‡https://huggingface.co/mixedbread-ai/mxbai-embed-large-v1 models as the first and second re-ranking Sentence-Transformer models, respectively.
+To re-rank the guided documents as in section 3.1.2, we select the ember-v1 222https://huggingface.co/llmrails/ember-v1 and mxbai-embed-large-v1 333https://huggingface.co/mixedbread-ai/mxbai-embed-large-v1 models as the first and second re-ranking Sentence-Transformer models, respectively.
 And we use OpenAI text-embedding-3-small for generating embeddings for keywords and answers in the filtering stage in GuideCQR.
@@
 We use KeyBERT (Grootendorst, 2020) for augmenting keywords.
-And we use RoBERTa-Base-Squad2 §§§https://huggingface.co/deepset/roberta-base-squad2 for generating the expected answer.
+And we use RoBERTa-Base-Squad2 444https://huggingface.co/deepset/roberta-base-squad2 for generating the expected answer.
 We use the most common query among the five candidate queries provided by the baseline Mao et al.
@@
 This means that the keyword span length is maximum number of keywords for each document.
-We conduct our experiments using an AMD EPYC 7313 CPU (3.0 GHz) paired with four NVIDIA RTX 4090 GPUs. We use Python 3.11.5 and PyTorch 2.3.1 for the software environment .
-Appendix B Query Analysis
-We present example queries reformulated by our method and by humans in Table 10.
-While our final query, , may not appear intuitive from a human perspective, it yields better search results during the retrieval stage.
-Appendix C Selecting FilterScore
+We use roberta-large-mnli 555https://huggingface.co/FacebookAI/roberta-large-mnli for measuring entailment score between augmented keywords and .
+In the query relevance analysis, we ensure a fair comparison by standardizing the number of documents in each set to 1,700 before conducting the experiment.
+We leverage Faiss index Johnson et al.
+(2019) for ANCE retriever.
+We conduct our experiments using an AMD EPYC 7313 CPU (3.0 GHz) paired with four NVIDIA RTX 4090 GPUs.
+We use Python 3.11.5 and PyTorch 2.3.1 for the software environment.
+Selecting FilterScore
 We experiment with various settings to find the optimal hyperparameter with the validation set for the proposed method.
-Before filtering, we use top-4 span 15 keywords and top-10 answers for CAsT-19, and top-5 span 5 keywords and top-10 answers for CAsT-20 and top-1 span 10 keywords and top-10 answer for QReCC.
+The number of documents for keywords and expected answers and optimal keyword span length varies by dataset; these values were chosen to improve retrieval results and enhance query quality for the retriever.
+Therefore, before filtering, we use top-4 span 15 keywords and top-10 answers for CAsT-19, and top-5 span 5 keywords and top-10 answers for CAsT-20 and top-1 span 10 keywords and top-10 answers for QReCC.
 And GuideCQR uses these keywords and answers as our final query set.
@@
 We achieve our final results using a of (1, 1.9) for each keyword and answer in CAsT-19, (0.1, 1.95) in CAsT-20, and (0.5, 9) in QReCC.
-Appendix D Process of generating of CAsT
-We generate of CAsT datasets by utilizing the LLM4CS REW GitHub code which employs GPT-3.5-turbo-16k for prompting Mao et al.
+| CAsT-19 | CAsT-20 | QReCC | Avg | |||||||
+| Methods | MRR | NDCG@3 | R@10 | MRR | NDCG@3 | R@10 | MRR | NDCG@3 | R@10 | |
+| RawQuery | 31.9 | 13.1 | 3.8 | 9.9 | 6.8 | 2.5 | 6.5 | 5.5 | 11.1 | 10.1 |
+| 59.9 | 29.5 | 10.3 | 30.5 | 19.0 | 10.7 | 46.9 | 55.0 | 65.5 | 36.4 | |
+| LLM4CS | 67.5 | 41.9 | 11.5 | 53.7 | 38.8 | 20.5 | 47.8 | 45.0 | 69.1 | 44.0 |
+| GuideCQR (ours) | 75.4 | 49.7 | 14.0 | 49.1 | 34.4 | 16.4 | 51.2 | 58.6 | 69.0 | 46.4 |
+| Human Rewrite | 61.3 | 30.9 | 10.9 | 37.7 | 23.9 | 14.2 | 39.7 | 36.2 | 62.5 | 35.3 |
+Appendix B Result using Sparse Retrieval
+We provide the performance of GuideCQR using sparse retrieval.
+This result is not included in the main table because prior studies on the CAsT series have predominantly focused on dense retrieval.
+Consequently, retrieval results for baseline methods are unavailable and the rewritten queries are not provided either.
+Thus, we conduct experiments with the available resources.
+We perform retrieval using BM25 Robertson et al.
+(2009), and indexing is performed via Pyserini Lin et al.
 (2021).
-As illustrated in Table 11, the prompt consists of three parts: instruction, demonstration, tail instruction.
-The final prompt is a concatenation of these three components.
-Using this prompt, GuideCQR generate by refining with , resolving any coreferences or omissions.
-Appendix E Dialog history usage in GuideCQR
-In conversational search, utilizing dialog history is crucial.
-GuideCQR leverages the conversation history in two key ways.
-First, we use dialog history when generating , as it helps refine the current query by resolving any ambiguities, coreferences, or omissions from previous interactions.
-Additionally, we incorporate history in filtering augmented keywords and expected answers.
-By measuring the similarity with the history, we can select the most relevant keywords or answers.
-Through these methods, GuideCQR effectively integrates history, demonstrating its capability as a conversational search system.
-Appendix F Dataset Statistics
+For sparse retrieval in LLM4CS Mao et al.
+(2023), we concatenate the rewritten query and response of generated LLM4CS final query.
+As shown in Table 7, GuideCQR significantly enhances performance metrics compared to the across all datasets.
+GuideCQR achieves state-of-the-art performance in terms of the average score.
+Similar to earlier observations, the scores for CAsT-20 are lower due to the unique characteristics of the CAsT datasets.
+Appendix C Dataset Statistics
 CAsT
 Table 8 presents statistics for the CAsT datasets, including the number of conversations, queries, and documents.
-Both CAsT datasets share the same document collection and provide passage-level relevance judgments, as well as human rewrites for each turn.
 | CAsT-19 | CAsT-20 | |
@@
 | # Questions | 0 | 0 | 748 |
+Appendix D Performance among Keyword Span Length
+We also investigate the impact of varying keyword spans in augmenting the keyword stage and present the result in Table 10.
+We observe that increasing the keyword span generally enhances performance to a certain extent.
+However, we find that too many signals can degrade performance due to inclusion of the redundant information in the query.
+| CAsT-19 | CAsT-20 | |||
+| MRR | NDCG@3 | MRR | NDCG@3 | |
+| 72.2 | 45.2 | 53.4 | 37.2 | |
+| 5 span | 76.9 | 49.9 | 57.7 | 41.5 |
+| 10 span | 78.8 | 52.0 | 53.7 | 39.8 |
+| 15 span | 81.0 | 52.2 | 54.1 | 40.0 |
+| 20 span | 80.1 | 52.0 | 54.6 | 39.8 |
+Appendix E Performance for removing duplicated keywords
+We conduct experiments by removing duplicate keywords, as non-duplicated keywords may lead to improved retrieval performance.
+Table 11 shows that duplicated keywords performed slightly better than unique ones, so we keep them duplicated.
+| MRR | NDCG@3 | |
+| Original | 81.0 | 52.2 |
+| Duplicate_removed | 78.3 | 49.9 |
+Appendix F Process of generating of CAsT
+We generate of CAsT datasets by utilizing the LLM4CS REW GitHub code which employs GPT-3.5-turbo-16k for prompting Mao et al.
+(2023).
+As illustrated in Table 13, the prompt consists of three parts: instruction, demonstration, tail instruction.
+The final prompt is a concatenation of these three components.
+Using this prompt, GuideCQR generate by refining with , resolving any coreferences or omissions.
+Appendix G Dialog history usage in GuideCQR
+In conversational search, utilizing dialog history is crucial.
+GuideCQR leverages the conversation history in two key ways.
+First, we use dialog history when generating , as it helps refine the current query by resolving any ambiguities, coreferences, or omissions from previous interactions.
+Additionally, we incorporate history in filtering augmented keywords and expected answers.
+By measuring the similarity with the history, we can select the most relevant keywords or answers.
+Through these methods, GuideCQR effectively integrates history, demonstrating its capability as a conversational search system.
+Appendix H Query Sample
+We present example queries reformulated by our method and by humans in Table 12.
+While our final query, , may not appear intuitive from a human perspective, it yields better search results during the retrieval stage.
+Appendix I Samples of Positive Queries on Datasets
+We provide positive query sample of each dataset.
+For CAsT, if the top-ranked (first position) retrieved document has a relevance score of 1 or higher, we classify it as a positive case.
+For the QReCC dataset, we refer to a retrieval result as a positive case if the gold passage is included in the top-1 result.
+Table 14, Table 15 and Table 16 shows positive query sample of CAsT-19, CAsT-20, and QReCC dataset.
 | RawQuery | What are its other competitors? |
@@
 |
+| CAsT-19 Dataset | |
+| Previous Turns |
+Query: What is throat cancer?
+Response: “” Query: Is it treatable? Response: “” Query: Tell me about lung cancer. Response: “” |
+| Original Query | What are its symptoms? |
+| Human Rewrite | What are lung cancer’s symptoms? |
+| LLM | What are the symptoms of lung cancer? |
+| GuideCQR Query | What are the symptoms of lung cancer?
+signs symptoms lung cancer symptoms lung cancer signs symptoms lung cancer cough signs symptoms lung cancer cough (hemoptysis), breath, wheezing, hoarseness coughing; pain chest, coughing |
+| Gold Passage (4) | Signs and symptoms of lung cancer typically occur only when the disease is advanced.
+Signs and symptoms of lung cancer may include: A new cough that doesn’t go away.
+Changes in a chronic cough or smoker’s cough.
+Coughing up blood, even a small amount.
+Shortness of breath.
+Chest pain.
+|
+| CAsT-20 Dataset | |
+| Previous Turns |
+Query: How do you know when your garage door opener is going bad?
+Response: Light Socket or Logic Board.
+If the light is still not coming on, it could be the light socket or the logic board.
+To determine which one it is, follow the steps below.
+Unplug the garage door opener, then immediately plug the opener back into the electrical outlet.
+Listen for the light relay on the logic board to click.
+If you hear a click, replace the light socket.
+If you do not hear a click, replace the logic board.
+There are four causes for the lights on a garage door opener not to come on: 1.
+Light bulb.
+2.
+Contacts in the light socket.
+3.
+Light socket.
+Query: Now it stopped working.
+Why?
+Response: Garage Door Opener Problems.
+So, when the garage door opener decides to take a day off, it can leave you stuck outside, probably during a rain or snowstorm.
+Though they may seem complicated, there really are several things most homeowners can do to diagnose and repair opener failures.
+And, if you are careful not to damage the door or the seal on the bottom of the door, use a flat shovel or similar tool to chip away at the ice.
+Once you get the door open, clear any water, ice or snow from the spot on the garage floor where the door rests when closed.
+|
+| Original Query | How much does it cost for someone to fix it? |
+| Human Rewrite | How much does it cost for someone to repair a garage door opener? |
+| LLM | How much does it typically cost to have a professional fix a garage door opener issue? |
+| GuideCQR Query | How much does it typically cost to have a professional fix a garage door opener issue?
+garage door opener repair cost garage door opener repair cost cost repair broken garage door opener garage door opener repair cost garage door opener repair cost fix garage door opener |
+| Gold Passage (4) | On average, homeowners report paying $210 to have a garage door opener repaired by a handyman.
+Most homeowners pay between $170 and $250 for the cost of such a repair project.
+The minimum reported cost to repair a broken garage door opener is $50, while the most costly repair may amount to $350.
+If an electrician is needed to repair faulty wiring, the cost of the repair will increase.
+The minimum reported cost to repair a broken garage door opener is $50, while the most costly repair may amount to $350.
+If an electrician is needed to repair faulty wiring, the cost of the repair will increase.
+|
+| QReCC Dataset | |
+| Previous Turns |
+Query: What are the main breeds of goat?,
+Answer: Abaza…Zhongwei |
+| Original Query | Tell me about boer goats. |
+| Human Rewrite | Tell me about boer goats. |
+| LLM | What can you tell me about boer goats in relation to the main breeds of goats? |
+| GuideCQR Query | What can you tell me about boer goats in relation to the main breeds of goats? goat boer breed south developed |
+| Gold Answer | The Boer goat is a breed of goat that was developed in South Africa in the early 1900s for meat production.
+Their name is derived from the Afrikaans (Dutch) word boer, meaning farmer.
+|
+| Gold Passage | Boer goat - Wikipedia CentralNotice Boer goat From Wikipedia, the free encyclopedia Jump to navigation Jump to search A commercial Boer goat buck Commercial Boer goat buck Boer goat, Doe The Boer goat is a breed of goat that was developed in South Africa in the early 1900s and is a popular breed for meat production.
+Their name is derived from the Afrikaans ( Dutch ) word boer , meaning farmer.
+Contents 1 Origins and characteristics 2 Doe 3 Crossbreeding 4 Show goats 5 References 6 External links Origins and characteristics [ edit ] The Boer goat was probably bred from the indigenous South African goats kept by the Namaqua , San , and Fooku tribes, with some crossing of Indian and European bloodlines being possible.
+They were selected for meat rather than milk production; due to selective breeding and improvement, the Boer goat has a fast growth rate and excellent carcass qualities, making it one of the most popular breeds of meat goat in the world.
+Boer goats have a high resistance to disease and adapt well to hot, dry semideserts.
+United States production is centered in west-central Texas , particularly in and around San Angelo and Menard .
+The original US breeding stock came from herds located in New Zealand.
+Only later were they imported directly from Africa Boer goats commonly have white bodies and distinctive brown heads.
+Some Boer goats can be completely brown or white or paint, which means large spots of a different color are on their bodies.
+Like the Nubian goat , they possess long, pendulous ears.
+They are noted for being docile, fast-growing, and having high fertility rates.
+Does are reported to have superior mothering skills as compared to other breeds.
+|